# Stylized facts in Web3

### v2

@@ 4.7. Token Statistics and Characteristics
1+24 Else Slope Intercept + - Ini Fin R L
SPY TradFi 1998/7/1 $\sim$ 2023/12/5 2 -0.86 3.9 NaN NaN 0.52 -8.4 0.048 0.34 2.5 2.5 -1.5
EUR TradFi 2002/10/21 $\sim$ 2023/12/5 2 -0.82 3.8 NaN NaN 0.1 0.3 0.091 0.41 2.7 2.7 -1.8
Binance BTC PoW+L1 2021/9/23 $\sim$ 2023/10/18 $<$1 -1.3 1.6 -0.26 -1.3 0.4 -1.6 0.13 0.99 2.4 2.5 -2.3
Binance XRP PoS+L1 2020/12/31 $\sim$ 2024/1/31 1 -4 1.7 -0.22 -1.2 0.82 1.1 0.12 0.88 2.3 2.4 -3
Binance ETH PoS+L1 2017/8/17 $\sim$ 2023/10/18 1 -9.1 2.5 -0.22 -1.2 1.5 -6.3 0.21 0.23 2.6 2.5 -1.8
$\text{arb}_{30bp}$ DAI L1 2021/9/23 $\sim$ 2023/10/18 12 1.5 1.5 -0.21 -1.2 0.2 -3 0.075 0.75 2.5 2.5 -1.5
$\text{arb}_{30bp}$ USDC L1 2021/9/23 $\sim$ 2023/10/18 5 0.12 1.5 -0.19 -1.1 0.18 -2.9 0.081 0.73 2.6 2.5 -1.5
$\text{arb}_{5bp}$ USDC L1 2021/9/23 $\sim$ 2023/10/18 <1 -1.1 1.6 -0.2 -1.1 0.27 -2.9 0.072 0.6 2.5 2.5 -1.5
$\text{arb}_{5bp}$ USDT L1 2021/9/23 $\sim$ 2023/10/18 <1 -1.7 1.5 -0.22 -1.1 0.41 -3.3 0.15 0.6 2.5 2.4 -1.7
Coinbase AAVE Saving+Defi 2020/12/15 $\sim$ 2023/11/6 $<$1 -2.7 1.7 -0.15 -1.1 1.2 -1.5 0.063 -0.024 2.8 2.9 -1.6
Coinbase ADA PoS+L1 2021/3/18 $\sim$ 2023/11/6 2 -9.1 1.6 -0.19 -1.3 1.5 -3 0.043 0.07 2.6 2.6 -2
Coinbase ALGO PoS+L1 2019/8/15 $\sim$ 2023/11/6 3 -8.5 2 -0.19 -1.3 0.64 -1.7 0.13 0.58 2.9 3 -1.9
Coinbase ATOM PoS+L1 2020/1/14 $\sim$ 2023/11/6 $<$1 -7.4 1.9 -0.18 -0.98 1.2 -2.4 0.08 0.46 2.7 2.9 -2.3
Coinbase AVAX PoS+L1 2021/9/30 $\sim$ 2023/11/6 4 -5.8 1.5 -0.2 -1 0.22 -2.2 0.14 0.41 2.8 2.8 -2.1
Coinbase AXS Gaming 2021/8/12 $\sim$ 2023/11/6 6 -5.1 1.4 -0.24 -1.2 0.24 0.38 0.15 0.57 2.2 2.7 -2
Coinbase BAL Exchange+Defi 2020/10/6 $\sim$ 2023/11/6 7 -3.7 1.7 -0.16 -1.4 0.97 -2.1 0.033 -0.075 3.2 2.9 -1.8
Coinbase BCH PoW+L1 2017/12/20 $\sim$ 2023/11/6 $<$1 -7.9 2.6 -0.27 -0.87 0.91 -2.9 0.21 0.74 2.4 2.5 -2.7
Coinbase BTC PoW+L1 2016/1/1 $\sim$ 2023/11/6 $<$1 -8.3 2.8 -0.2 -1.1 1.3 -6.2 0.23 0.48 2.5 2.4 -2.3
Coinbase CRV Exchange+Defi 2021/3/25 $\sim$ 2023/11/6 1 -2.4 1.5 -0.19 -0.92 0.68 -2 0.085 0.37 2.8 2.8 -2.1
Coinbase CVX Stackingt+Defi 2022/9/8 $\sim$ 2023/11/6 18 -2.5 0.99 -0.33 -1.5 -0.55 1 0.16 0.87 2.5 2.5 -1.9
Coinbase DOGE PoS+Memecoin 2021/6/3 $\sim$ 2023/11/6 2 -5.3 1.5 -0.33 -1.4 0.4 0.36 0.15 0.64 2.3 2.6 -2.1
Coinbase DOT PoS+L1 2021/6/16 $\sim$ 2023/11/6 1 -3.5 1.4 -0.23 -1 0.33 -1.9 0.14 0.48 2.7 2.8 -1.9
Coinbase ETH PoS+L1 2016/5/18 $\sim$ 2023/11/6 1 -7.6 2.9 -0.24 -1.1 2.7 -5.6 0.21 0.56 2.5 2.5 -1.8
Coinbase GNO PoS+L1 2022/7/21 $\sim$ 2023/11/6 24 -11 0.99 -0.35 -0.96 0.041 -0.063 0.053 0.77 3.3 3.3 -1.7
Coinbase INJ Exchange+Defi 2022/9/20 $\sim$ 2023/11/6 3 -4.1 1.2 -0.19 -1.2 0.19 1.9 0.043 -0.55 2.7 2.7 -1.6
Coinbase LINK Oracle 2019/6/27 $\sim$ 2023/11/6 $<$1 -9.1 2.1 -0.23 -1.5 1.5 -3 0.16 0.67 2.9 2.7 -2.2
Coinbase LTC PoW+L1 2016/8/17 $\sim$ 2023/11/6 4 -11 2.7 -0.25 -0.8 1.7 -1 0.2 0.64 2.5 2.6 -1.6
Coinbase MATIC L2 2021/3/11 $\sim$ 2023/11/6 $<$1 -6.7 1.7 -0.15 -1.2 2.1 0.31 0.097 0.42 2.5 2.8 -2.2
Coinbase OP L2 2022/6/1 $\sim$ 2023/11/6 4 -0.15 1.3 -0.19 -0.82 -0.41 -1 0.15 1.4 2.5 2.6 -1.8
Coinbase SAND Gaming 2022/5/26 $\sim$ 2023/11/6 2 -0.91 1.2 -0.21 -1.1 0.26 0.13 0.077 0.34 2.7 2.6 -1.8
Coinbase SHIB PoS+Memecoin 2022/1/1 $\sim$ 2023/11/6 8 -7.4 1.4 -0.34 -1.2 0.29 0.015 0.2 0.71 2.3 2.5 -2.6
Coinbase SNX Exchange+Defi 2020/12/15 $\sim$ 2023/11/6 2 -5.9 1.5 -0.22 -1.2 0.7 0.55 0.12 -0.05 2.8 3 -2.3
Coinbase SOL PoS+L1 2021/6/17 $\sim$ 2023/11/6 2 -4.1 1.8 -0.33 -1.2 0.49 -2.2 0.17 0.76 2.6 2.9 -1.8
Coinbase UNI Exchange+Defi 2020/9/17 $\sim$ 2023/11/6 1 -5.1 1.6 -0.19 -1.2 3 0.62 0.023 -0.032 2.7 2.8 -1.7
$\text{m}_{30bp}$ DAI L1 2021/9/23 $\sim$ 2023/10/18 19 -0.21 1.5 -0.19 -1.2 0.18 -3.1 0.094 0.69 2.5 2.5 -1.6
$\text{m}_{30bp}$ USDC L1 2021/9/23 $\sim$ 2023/10/18 8 -0.46 1.5 -0.18 -1.2 0.15 -3.2 0.1 0.68 2.4 2.5 -1.5
$\text{m}_{30bp}$ USDT L1 2021/9/23 $\sim$ 2023/10/18 4 -0.64 1.6 -0.18 -1.2 0.14 -3.1 0.089 0.65 2.5 2.5 -1.5
$\text{m}_{5bp}$ USDC L1 2021/9/23 $\sim$ 2023/10/18 6 -1.8 1.6 -0.18 -1.2 0.17 -3 0.095 0.51 2.4 2.5 -1.5
@@ 5.5.1. Arbitrage Mechanics in DEX
$Z_{\text{opt}}=\begin{cases}\gamma^{-1}S&\text{if }Z>\gamma^{-1}S\\ Z&\text{if }Z\in[\gamma S,\gamma^{-1}S]\\ \gamma S&\text{if }Z<\gamma S\end{cases}$
The interval $[\gamma S,\gamma^{-1}S]$ is termed the "no-arbitrage region," where arbitrageurs cannot profit due to trading fees. We refer the interested reader to [TW20, AC20, MMR23] for further details.
@@ 5.5.2. Empirical Observations
Figure 16 suggests that CEX leads DEX in price discovery. The price discovery appears to happen on CEX, and only when it is profitable does the price on a DEX change to match the new CEX price [AC20, MMR23, NTYY24]. We can further formalize this observation by examining the correlation between lagged ETH returns on a CEX and on a DEX.
@@ 5.5.3. Price Discovery and Lead-Lag Relationships
$\frac{\mathbb{E}\left[R_{B}(t)R_{p}(t+k)\right]}{\sigma_{{R_{B}}}\sigma_{{R_{p}}}}$
@@ 5.5.4. Implications of Fee Structures
Furthermore, the differences in behavior between 30bp and 5bp pools highlight how fee structures can significantly impact trading dynamics and market efficiency. The higher fees in the 30bp pool appear to create a more stable price environment, while the lower fees in the 5bp pool lead to more frequent price adjustments, potentially offering more opportunities for nimble traders [MMR23].
@@ 6. Acknowledgement
The authors thank Wun-Cing Liou and Min-Ren Guan for their invaluable assistance in collecting blockchain data. We also thank Wei-Lun Tao for his crucial contribution to cleaning traditional finance data.
ACS acknowledges the generous support and hospitality provided by the National Center for Theoretical Sciences, Mathematics Division, at the National Taiwan University, and the Department of Mathematics at the National Tsing Hua University during the summer of 2024.

### v3

@@ 4.7. Token Statistics and Characteristics
1+24 Else Slope Int. + - Ini Fin R L
SPY TradFi 1998/7/1–2023/12/5 2 -0.86 3.9 NaN NaN 0.52 -8.4 0.048 0.34 2.5 2.5 -1.5
EUR TradFi 2002/10/21–2023/12/5 2 -0.82 3.8 NaN NaN 0.1 0.3 0.091 0.41 2.7 2.7 -1.8
Binance BTC PoW+L1 2021/9/23–2023/10/18 $<$1 -1.3 1.6 -0.26 -1.3 0.4 -1.6 0.13 0.99 2.4 2.5 -2.3
Binance XRP PoS+L1 2020/12/31–2024/1/31 1 -4 1.7 -0.22 -1.2 0.82 1.1 0.12 0.88 2.3 2.4 -3
Binance ETH PoS+L1 2017/8/17–2023/10/18 1 -9.1 2.5 -0.22 -1.2 1.5 -6.3 0.21 0.23 2.6 2.5 -1.8
$\text{arb}_{30bp}$ DAI L1 2021/9/23–2023/10/18 12 1.5 1.5 -0.21 -1.2 0.2 -3 0.075 0.75 2.5 2.5 -1.5
$\text{arb}_{30bp}$ USDC L1 2021/9/23–2023/10/18 5 0.12 1.5 -0.19 -1.1 0.18 -2.9 0.081 0.73 2.6 2.5 -1.5
$\text{arb}_{5bp}$ USDC L1 2021/9/23–2023/10/18 $<$1 -1.1 1.6 -0.2 -1.1 0.27 -2.9 0.072 0.6 2.5 2.5 -1.5
$\text{arb}_{5bp}$ USDT L1 2021/9/23–2023/10/18 $<$1 -1.7 1.5 -0.22 -1.1 0.41 -3.3 0.15 0.6 2.5 2.4 -1.7
Coinbase AAVE Saving+Defi 2020/12/15–2023/11/6 $<$1 -2.7 1.7 -0.15 -1.1 1.2 -1.5 0.063 -0.024 2.8 2.9 -1.6
Coinbase ADA PoS+L1 2021/3/18–2023/11/6 2 -9.1 1.6 -0.19 -1.3 1.5 -3 0.043 0.07 2.6 2.6 -2
Coinbase ALGO PoS+L1 2019/8/15–2023/11/6 3 -8.5 2 -0.19 -1.3 0.64 -1.7 0.13 0.58 2.9 3 -1.9
Coinbase ATOM PoS+L1 2020/1/14–2023/11/6 $<$1 -7.4 1.9 -0.18 -0.98 1.2 -2.4 0.08 0.46 2.7 2.9 -2.3
Coinbase AVAX PoS+L1 2021/9/30–2023/11/6 4 -5.8 1.5 -0.2 -1 0.22 -2.2 0.14 0.41 2.8 2.8 -2.1
Coinbase AXS Gaming 2021/8/12–2023/11/6 6 -5.1 1.4 -0.24 -1.2 0.24 0.38 0.15 0.57 2.2 2.7 -2
Coinbase BAL Exchange+Defi 2020/10/6–2023/11/6 7 -3.7 1.7 -0.16 -1.4 0.97 -2.1 0.033 -0.075 3.2 2.9 -1.8
Coinbase BCH PoW+L1 2017/12/20–2023/11/6 $<$1 -7.9 2.6 -0.27 -0.87 0.91 -2.9 0.21 0.74 2.4 2.5 -2.7
Coinbase BTC PoW+L1 2016/1/1–2023/11/6 $<$1 -8.3 2.8 -0.2 -1.1 1.3 -6.2 0.23 0.48 2.5 2.4 -2.3
Coinbase CRV Exchange+Defi 2021/3/25–2023/11/6 1 -2.4 1.5 -0.19 -0.92 0.68 -2 0.085 0.37 2.8 2.8 -2.1
Coinbase CVX Stacking+Defi 2022/9/8–2023/11/6 18 -2.5 0.99 -0.33 -1.5 -0.55 1 0.16 0.87 2.5 2.5 -1.9
Coinbase DOGE PoS+Memecoin 2021/6/3–2023/11/6 2 -5.3 1.5 -0.33 -1.4 0.4 0.36 0.15 0.64 2.3 2.6 -2.1
Coinbase DOT PoS+L1 2021/6/16–2023/11/6 1 -3.5 1.4 -0.23 -1 0.33 -1.9 0.14 0.48 2.7 2.8 -1.9
Coinbase ETH PoS+L1 2016/5/18–2023/11/6 1 -7.6 2.9 -0.24 -1.1 2.7 -5.6 0.21 0.56 2.5 2.5 -1.8
Coinbase GNO PoS+L1 2022/7/21–2023/11/6 24 -11 0.99 -0.35 -0.96 0.041 -0.063 0.053 0.77 3.3 3.3 -1.7
Coinbase INJ Exchange+Defi 2022/9/20–2023/11/6 3 -4.1 1.2 -0.19 -1.2 0.19 1.9 0.043 -0.55 2.7 2.7 -1.6
Coinbase LINK Oracle 2019/6/27–2023/11/6 $<$1 -9.1 2.1 -0.23 -1.5 1.5 -3 0.16 0.67 2.9 2.7 -2.2
Coinbase LTC PoW+L1 2016/8/17–2023/11/6 4 -11 2.7 -0.25 -0.8 1.7 -1 0.2 0.64 2.5 2.6 -1.6
Coinbase MATIC L2 2021/3/11–2023/11/6 $<$1 -6.7 1.7 -0.15 -1.2 2.1 0.31 0.097 0.42 2.5 2.8 -2.2
Coinbase OP L2 2022/6/1–2023/11/6 4 -0.15 1.3 -0.19 -0.82 -0.41 -1 0.15 1.4 2.5 2.6 -1.8
Coinbase SAND Gaming 2022/5/26–2023/11/6 2 -0.91 1.2 -0.21 -1.1 0.26 0.13 0.077 0.34 2.7 2.6 -1.8
Coinbase SHIB PoS+Memecoin 2022/1/1–2023/11/6 8 -7.4 1.4 -0.34 -1.2 0.29 0.015 0.2 0.71 2.3 2.5 -2.6
Coinbase SNX Exchange+Defi 2020/12/15–2023/11/6 2 -5.9 1.5 -0.22 -1.2 0.7 0.55 0.12 -0.05 2.8 3 -2.3
Coinbase SOL PoS+L1 2021/6/17–2023/11/6 2 -4.1 1.8 -0.33 -1.2 0.49 -2.2 0.17 0.76 2.6 2.9 -1.8
Coinbase UNI Exchange+Defi 2020/9/17–2023/11/6 1 -5.1 1.6 -0.19 -1.2 3 0.62 0.023 -0.032 2.7 2.8 -1.7
$\text{m}_{30bp}$ DAI L1 2021/9/23–2023/10/18 19 -0.21 1.5 -0.19 -1.2 0.18 -3.1 0.094 0.69 2.5 2.5 -1.6
$\text{m}_{30bp}$ USDC L1 2021/9/23–2023/10/18 8 -0.46 1.5 -0.18 -1.2 0.15 -3.2 0.1 0.68 2.4 2.5 -1.5
$\text{m}_{30bp}$ USDT L1 2021/9/23–2023/10/18 4 -0.64 1.6 -0.18 -1.2 0.14 -3.1 0.089 0.65 2.5 2.5 -1.5
$\text{m}_{5bp}$ USDC L1 2021/9/23–2023/10/18 6 -1.8 1.6 -0.18 -1.2 0.17 -3 0.095 0.51 2.4 2.5 -1.5
@@ 5.2. Quantitative Differences Among Tokens
5.3. Time Stability of Stylized Facts
The statistics studied in this article can exhibit significant temporal variation. We examine this variation through BTC and ETH, focusing on two distinct temporal frameworks: market regimes and the May 2020 halving event.
5.3.1. Market Regime Analysis
To deepen our understanding of token market dynamics, we compare the stylized facts of BTC and ETH in two market regimes after 2020. This period is particularly interesting as it encompasses two contrasting market phases: a significant bull market from March 2020 to April 2021 and a pronounced bear market from December 2021 to January 2023.
Figure 16 illustrates these distinct market phases through the scaled price movements of both tokens. The contrast between the bull and bear markets provides an excellent natural experiment for examining how market conditions affect the statistical properties of these assets.
Figure 16. Scaled price of ETH and BTC starting from June 2017. The graph illustrates a bull market and a bear market after 2020.
Our empirical analysis reveals several key findings. Table 2 shows that certain fundamental properties remain stable across market regimes. The unconditional return distribution, measured by the tail of the cumulative distribution function (CDF Tail), and volatility clustering (ACF Abs(R)) demonstrate remarkable consistency between regimes. Both ETH and BTC exhibit similar patterns in these metrics. While these findings broadly align with Table 1, we observe an important distinction in the absolute return autocorrelation. The full dataset shows a faster decay with a slope of approximately $-0.22$ (Table 1), compared to $-0.13$ in our regime analysis (Table 2). Section 5.3.2 attributes this difference to structural changes around the May 2020 halving event.
Token Regime History Zeros(%) Avg ACF ACF Abs(R) Avg Lev TRA CDF Tail JB
1+24 Else Slope Int. + - Ini Fin R L
BTC Bear 2021/12/1–2023/1/1 $<$1 -3.1 2.9 -0.16 -1.2 0.37 -2.0 0.13 0.95 2.3 2.3
BTC Bull 2020/3/1–2021/4/1 $<$1 -7.9 2.8 -0.12 -1.1 0.73 -3.4 0.26 0.13 2.4 2.2
ETH Bear 2021/12/1–2023/1/1 $<$1 -2.6 3.0 -0.14 -1.2 0.41 -2.6 0.11 0.63 2.3 2.3
ETH Bull 2020/3/1–2021/4/1 $<$1 -7.5 2.9 -0.12 -1.1 0.6 -3.7 0.22 -0.23 2.5 2.4
Table 2. Statistics in Bull and Bear markets
Other stylized facts exhibit notable regime-dependent behavior. Return autocorrelation patterns differ significantly from the full sample statistics (Table 1). During the bear market, both assets show insignificant autocorrelation at three standard errors, with "Avg ACF" values generally within ±$\pm 3$ (BTC slightly lower at $-3.1$). However, the first and 24th-hour return autocorrelations fall well below $3$ in the bear market period. This pattern requires careful interpretation, as it might reflect market evolution rather than regime effects. Indeed, the more recent bear market period would naturally exhibit greater efficiency, and Section 5.3.2 suggests this evolutionary interpretation.
The leverage effect presents distinct patterns across regimes. Table 2 shows negative cross-correlation (column "Ave Lev,-") between current returns and future absolute returns in both periods, but with varying intensity. The effect is weaker during the bear market (values above $-3$) compared to the bull market (values below $-3$: $-3.4$ for BTC, $-3.7$ for ETH). This asymmetry aligns with the characteristic volatility patterns of these regimes: the high volatility of bear markets may dampen the leverage effect, while the lower volatility during bull markets may amplify it.
Perhaps the most striking finding emerges from the Time Reversal Asymmetry (TRA) analysis. Table 2 reveals a clear dichotomy: the TRA effect is notably absent during the bull market period but manifests strongly during the bear market phase.
5.3.2. Halving Event Analysis
To further examine the temporal stability of stylized facts, we analyze ETH and BTC across a significant structural break: the May 2020 Bitcoin halving event, when mining rewards decreased from 12.5 to 6.25 BTC. We select these tokens due to their extensive historical data and split our analysis into pre- and post-May 2020 periods. Our dataset encompasses this single halving event.
Token Halving History Zeros(%) Avg ACF ACF Abs(R) Avg Lev TRA CDF Tail JB
1+24 Else Slope Int. + - Ini Fin R L
BTC Before 2017/8/17–2020/4/30 $<$1 -9.0 1.7 -0.19 -1.2 0.7 -4.2 0.28 0.56 2.4 2.4
BTC After 2020/5/1–2023/10/19 $<$1 -3.6 1.9 -0.16 -1.1 1.3 -3.0 0.13 0.38 2.7 2.6
ETH Before 2017/8/17–2020/4/30 $<$1 -8.5 1.7 -0.29 -1.2 0.9 -4.4 0.29 0.41 2.5 2.4
ETH After 2020/5/1–2023/10/19 $<$1 -2.9 2.0 -0.14 -1.1 1.6 -4.1 0.09 0.05 2.8 2.5
Table 3. Statistics before and after May 2020
Table 3 reveals several persistent statistical properties across this structural break. The unconditional probability density maintains heavy tails with tail coefficients showing minimal temporal variation. Similarly, the leverage effect remains consistently present, with comparable magnitudes in both periods.
Time Reversal Asymmetry exhibits asset-specific temporal variation. While BTC maintains significant TRA effects throughout both periods, ETH shows a marked change, with TRA disappearing in the post-2020 period. This finding aligns with our earlier observations in Table 1 and Table 2, reinforcing that TRA’s presence depends on both the specific token and the time period under consideration.
Volatility clustering persists across both periods but shows distinct characteristics. The autocorrelation function slopes indicate stronger persistence before May 2020 ($-0.19$ for BTC, $-0.29$ for ETH) compared to the post-halving period ($-0.16$ for BTC, $-0.14$ for ETH). This pattern aligns with our findings in Section 5.3.1 regarding the bull and bear market regimes, which fall within the second half of our sample. Moreover, when comparing these results with the full sample statistics in Table 1, we observe that the first half of the data (pre-May 2020) dominates the overall sample characteristics.
The log returns’ autocorrelation provides particularly instructive insights about market evolution. This metric, often considered the most fundamental measure of market efficiency, shows substantially stronger negative first-lag autocorrelation in the first half of the sample. This temporal pattern suggests a systematic increase in market efficiency following the halving event.
@@ 5.6.1. Arbitrage Mechanics in DEX
$Z_{\text{opt}}=\begin{cases}\gamma^{-1}S&\text{if }Z>\gamma^{-1}S\\ Z&\text{if }Z\in[\gamma S,\gamma^{-1}S]\\ \gamma S&\text{if }Z<\gamma S\end{cases}$ (5)
The interval $[\gamma S,\gamma^{-1}S]$ is termed the "no-arbitrage region," where arbitrageurs cannot profit due to trading fees. We refer the interested reader to [18, 1, 10] for further details.
@@ 5.6.2. Empirical Observations
Figure 18 suggests that CEX leads DEX in price discovery. The price discovery appears to happen on CEX, and only when it is profitable does the price on a DEX change to match the new CEX price [1, 10, 11]. We can further formalize this observation by examining the correlation between lagged ETH returns on a CEX and on a DEX.
@@ 5.6.3. Price Discovery and Lead-Lag Relationships
$\frac{\mathbb{E}\left[R_{B}(t)R_{p}(t+k)\right]}{\sigma_{{R_{B}}}\sigma_{{R_{p}}}}$ (6)
@@ 5.6.4. Implications of Fee Structures
Furthermore, the differences in behavior between 30bp and 5bp pools highlight how fee structures can significantly impact trading dynamics and market efficiency. The higher fees in the 30bp pool appear to create a more stable price environment, while the lower fees in the 5bp pool lead to more frequent price adjustments, potentially offering more opportunities for nimble traders [10].
@@ 6. Acknowledgement
The authors gratefully acknowledge Wun-Cing Liou and Min-Ren Guan for their essential contributions to blockchain data collection and Wei-Lun Tao for his meticulous work in cleaning traditional finance data.
ACS acknowledges the generous support and hospitality of the National Center for Theoretical Sciences, Mathematics Division, at the National Taiwan University, and the Department of Mathematics at the National Tsing Hua University during the summer of 2024.
This research was supported in part by NSTC grant 111-2115-M-007-014-MY3 (S.N. Tung).